# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

### v3

@@ V-A Experimental Setting
* re-implemented by [61]; \$ reference-based modification of a reference-less method; \$\$ reference-less modification of a reference-based method; $\dagger$ implemented by [76]; $\ddagger$ prompt-based modification of a reference-based method; $+$ results on the validation set computed by [84].
@@
TABLE III: Results on the CARPK dataset [44]. We report the performance in terms of MAE and RMSE obtained with some of the methodologies discussed in this survey over the CARPK benchmark. Following previous literature, CARPK is exploited for testing cross-dataset generalization over a specific object class – vehicles. The best results are in bold.
@@
V-B Reference-based Results
Table II shows the results achieved by the surveyed reference-based methods on the FSC-147 dataset [37]. We additionally report the results concerning two trivial baselines as in [37], i.e., always output the average object count for training images and always output the median count for the training images. The pioneering FamNet [37] laid the groundwork for this paradigm with an MAE and an RMSE of 22.08 and 99.54 on the test subset. Subsequent works progressively improve the performance, sometimes revising and adapting this foundation model. For instance, VCN [59], RCAC [60], and SPDCN [62] focused on the robustness of exemplar feature representations, improving generalization and highlighting the importance of robust training data diversity. Especially VCN and SPDCN reached an impressive MAE of 18.17 and 13.51, respectively. Differently, MACnet introduced segmentation masks as an alternative to bounding boxes for exemplar representation, demonstrating that finer object delineation can improve model robustness. The subsequent method that has marked a breakthrough is BMNet+, which was one of the first works that improved the matching technique between image and exemplar features, moving away from the naive idea of the fixed inner product between them. SAFECount [64] also followed the direction of improving the image-exemplars matching mechanism and achieved an MAE and RMSE of 14.32 and 85.54, respectively (again, on the test subset). Other methods that stood out are SSD [67], LOCA [65], and CACViT [71] which sensibly lowered both MAE and RMSE. The latter approach proposed a particularly interesting "extract-and-match" paradigm, which basically totally relies on the attention mechanism of Video Transformer [112]. Additionally, also ConCoNet is worth mentioning since it proposed a training strategy potentially applicable to all the other methods that brought the idea of leveraging negative exemplars to better distinguish between target and background regions. However, the model achieving state-of-the-art results on FSC-147 is DAVE [72]. By employing a detect-and-verify strategy, DAVE combines bounding-box detection with clustering-based verification to eliminate outliers effectively. On the leaderboard, DAVE demonstrates the lowest MAE and RMSE values to date (8.66 and 32.36, respectively, on the test subset), showcasing its dominance in the reference-based paradigm.
V-C Reference-less Results
In Table II, we report the results obtained by the surveyed reference-less approaches on the FSC-147 dataset [37]. The first reference-less approach, RepRPN-Counter [76], introduced the concept of repetitive region proposals and reached an MAE of 26.66 and an RMSE of 129.11 on the test subset. These error metrics are significantly higher than those obtained with reference-based approaches, as expected since reference-less techniques do not utilize bounding box exemplar annotations. Subsequent methods have significantly improved performance. CounTR [78], a transformer-based reference-less framework, leverages the self-similarity of image regions and achieves competitive results, with an MAE of 14.12 and an RMSE of 108.01 on the test subset. Another notable contribution is RCC [77]. Besides achieving competitive results (an MAE of 17.12 and an RMSE of 104.53), it employs a weakly supervised learning approach that does not need point-level annotations, thereby representing a particularly interesting approach in data scarcity scenarios. In the same direction, also GCNet [79] does not use dot labels and obtained comparable results to RCC. Overall, there is not a clear winner in this leaderboard. Indeed, on the one hand, the reference-less version of DAVE [72] achieves the best results in the validation split with an MAE and RMSE of 15.54 and 52.67 respectively; on the other hand, CounTR and GCNet shine in the test subset with an MAE of 14.12 and an RMSE of 102.89.
V-D Open-world Text-guided Results
In the last part of Table II, we show the results concerning the considered open-world text-guided approaches, which represent a recent and promising advancement in class-agnostic counting, driven by the progress of large multi-modal models. However, similar to the reference-less case, this flexibility comes at the expense of lower performance compared to reference-based techniques. The first milestone, ZSC [80], obtained an MAE and an RMSE of 22.09 and 115.17 on the test subset. Among the subsequent methods that steadily improved performance, CounTX [81], emerges as a particularly competitive approach, in some way adapting its counterpart CounTR [78] to this new setting. It achieves an impressive MAE of 15.73 and an RMSE of 106.88 on the test subset. However, the best results were contended for by the open-world text-guided variant of DAVE [72] and GroundingREC [84]. On the one hand, the first one excels in the RMSE metric, i.e., is more robust to outliers. On the other hand, the latter prevails on the MAE (10.06 and 10.12 on the validation and test subset, respectively).
V-E Cross-dataset Generalization
Table V-A summarizes the results on the CARPK dataset [44], focusing on cross-dataset generalization among the methods that reported their outcomes on this benchmark. Fine-tuned reference-based approaches demonstrate substantial performance gains, with CACViT [71] leading the pack, achieving an MAE of 4.91 and an RMSE of 6.49. This highlights the effectiveness of incorporating transformers for robust feature extraction and attention mechanisms. BMNet+ [61] and SAFECount [64] also deliver strong results, with MAEs of 5.76 and 5.33, respectively, leveraging advanced similarity modules and data-driven refinement strategies. Non-fine-tuned methods generally show higher error rates, yet recent advancements like LOCA [65], SATCount [68], CACViT [71], and SSD [67] deliver impressive performance, narrowing the gap fine-tuned approaches. Notably, the best-performing non-fine-tuned method is UPC [74], achieving an MAE of 7.84 and an RMSE of 9.74. In contrast, for open-world text-guided methods, VLCounter [83] stands out as the top performer, with an MAE of 6.46 and an RMSE of 8.68, even if UPC still achieved very competitive results (MAE of 7.62 and RMSE of 9.71).
V-F Textual Semantic Understanding
@@
Recently, [124] highlighted notable weaknesses in state-of-the-art open-world text-guided class-agnostic counting methods. Specifically, the authors argued that these models often fail to accurately identify which object class needs to be counted based on the textual description. Instead, they tend to count instances of the predominant class, disregarding the true intent of the prompt. To address these shortcomings, the authors introduced a novel benchmark, including a test suite specifically designed to evaluate these limitations in current evaluation systems. Table IV presents results on the FSC-147 dataset, focusing on their negative-label test. This test evaluates the model’s performance on single-class images by prompting it with textual descriptions referencing absent object classes. [124] introduced also two metrics for the assessment of this test: (i) the normalized mean of negative predictions (NMN), which is the absolute counting error computed by prompting the model with the negative classes, normalized by the ground truth of the positive class, and (ii) the positive class count nearness (PCCN) that mixes positive and negative class predictions, providing an overall quantitative assessment of strong failures in the model. More details can be found in [124]. DAVE [72] performs exceptionally well across nearly all metrics proposed in [124], standing out as the best-performing model in this challenging scenario. Although the authors noted that DAVE occasionally produces catastrophic errors due to sporadic outliers, it remains the top performer overall in this evaluation.
@@ VI Conclusion and Open Challenges
Class-agnostic counting (CAC) has become a pivotal field in object counting, addressing the limitations of traditional class-specific methods. Unlike conventional approaches that require extensive labeled data for each object category, CAC enables models to count objects across arbitrary categories, including those unseen during training. For the first time, this survey reviewed the evolution of CAC methodologies, categorizing them into three paradigms: reference-based, reference-less, and open-world text-guided approaches. Each paradigm brings distinct advantages and trade-offs, reflecting the diverse demands of real-world applications.
Reference-based methods, which rely on annotated exemplars, have established strong performance benchmarks by leveraging exemplar-guided mechanisms. These methods excel in precision and robustness, particularly in structured scenarios. However, their reliance on labeled exemplars limits scalability, making them less suitable for tasks requiring minimal human intervention. Reference-less methods address this limitation by automatically identifying dominant object classes in an image, eliminating the need for exemplars. While these approaches demonstrate greater flexibility and ease of use, their performance still lags behind exemplar-based methods, particularly in complex or cluttered scenes. Further refinements are needed to close this gap. The latest paradigm, open-world text-guided counting, introduces the ability to specify object classes through textual prompts, leveraging advancements in vision-language models. This approach offers unprecedented adaptability and aligns closely with the demands of dynamic and diverse environments. However, as with reference-less methods, this flexibility comes at the cost of reduced accuracy compared to exemplar-based approaches.
Through an in-depth analysis of state-of-the-art methods, including benchmarks such as FSC-147 [37] and CARPK [44], this survey highlights notable advancements and persistent challenges. Across paradigms, a clear ranking of models based on performance metrics does not always emerge. DAVE [72] leads the field in reference-based counting on FSC-147, achieving the lowest MAE (8.66) and RMSE (32.36). Its detect-and-verify strategy sets a benchmark for accuracy and robustness. Among reference-less models, instead, the FSC-147 leaderboard remains inconclusive. On the one hand, DAVE excels, achieving the best result on the validation split. On the other hand, CounTR [78] and GCNet [79] stand out in the test subset. The former showcases the effectiveness of its transformer-based design, while the latter is particularly noteworthy for its weakly supervised methodology, which does not rely on point-level annotations. Similarly, in open-world text-guided counting, FSC-147 lacks a clear winner. While DAVE often tops the leaderboard, demonstrating robustness to outliers, GroundingREC [84] achieves the lowest MAE in both the validation and test splits. Regarding generalization capabilities, the CARPK leaderboard indicates that CACViT [71], UPC [74], and VLCounter [83] are the top-performing approaches across different paradigms and experimental settings. Finally, an analysis of textual semantic understanding – applicable only to open-world text-guided methods – reveals that DAVE remains the best-performing approach.
@@
Open challenges still remain in class-agnostic counting and affect all paradigms. The first limitation to overcome is the lack of publicly available datasets. The gold standard, FSC-147, has several drawbacks. First, it contains a limited number of images and object classes. Furthermore, as pointed out in [124], almost all images contain only a single class of objects. Multi-class images would be a desirable feature for new datasets, as objects belonging to classes different from the one being counted could act as distractors for the counting approach. Finally, the provided textual descriptions are also insufficient, as they consist only of class names, whereas a desirable improvement would be to include more fine-grained class descriptions. The lack of diverse and large-scale datasets also leads to challenges in designing learning approaches suited for data-scarce scenarios. To date, only two methodologies –RCC [77] and GCNet [79] – operate with weak supervision, avoiding the use of dot annotations. This setting is particularly promising, as it could help mitigate the challenge of creating large-scale labeled datasets for the counting task. Other challenges are more specific to open-world text-guided methodologies, which seem to represent the future, especially considering the ubiquity of LMMs that we are witnessing. Specifically, they have been shown to not always correctly interpret the meaning of the prompt. Even though some models appear relatively robust, improvements are needed, and new architectural designs and learning approaches – such as contrastive learning over classes not present in the images – should be considered.

### v4

@@ IV Datasets
* re-implemented by [63]; \$ reference-based modification of a reference-less method; \$\$ reference-less modification of a reference-based method; $\dagger$ implemented by [79]; $\ddagger$ prompt-based modification of a reference-based method; ++ results on the validation set computed by [87]; ** trained on the FSC-147-D dataset [84]; ++ trained on the REC-8K dataset [87].
V Results and Discussion
In this section, we present the results of the evaluated methods, following the same taxonomy of previous sections—reference-based, reference-less, and open-world text-guided approaches. First, we describe the experimental setting and then provide a critical analysis of the outcomes across various points of discussion.
V-A Experimental Setting
Some techniques have been adapted to carry out more than one of these tasks (see also Sec. III); in these cases, we report the results for all of them. We considered the two most commonly used benchmarks: FSC-147 [32] and CARPK [43]. For FSC-147, in line with existing literature, we report results for both the validation and test subsets. Differently, the CARPK dataset is used to evaluate the generalization capabilities of the models in counting instances of a specific object class—vehicles. Previous works have explored two different settings: i) training models on FSC-147 without using any data from CARPK nor the car category in FSC-147, and ii) fine-tuning models using training data from CARPK, where a set of 12 exemplars is randomly sampled from the training set and used as exemplars across all training and test images.
V-B Reference-based Results
Table II shows the results achieved by the surveyed reference-based methods on FSC-147. We also report the results concerning two trivial baselines, as in [32], i.e., always output the average object count for training images and always output the median count for training images.
The pioneering FamNet [32] laid the groundwork for this paradigm with an MAE and an RMSE of 22.08 and 99.54 on the test subset. Subsequent works progressively improve the performance, sometimes revising and adapting this foundation model. For instance, VCN [61], RCAC [62], and SPDCN [64] focused on the robustness of exemplar feature representations, improving generalization, and highlighting the importance of robust training data diversity. Especially, VCN and SPDCN reached an impressive MAE of 18.17 and 13.51, respectively. Differently, MACnet introduced segmentation masks as an alternative to bounding boxes for exemplar representation, demonstrating that finer object delineation can improve model robustness. The subsequent method that has marked a breakthrough is BMNet+, which was one of the first works that improved the matching technique between image and exemplar features, moving away from the naive idea of the fixed inner product between them. SAFECount [66] also followed the direction of improving the image-exemplars matching mechanism and achieved an MAE and RMSE of 14.32 and 85.54, respectively (again, on the test subset). Other methods that stood out are SSD [69], LOCA [67], and CACViT [73], which sensibly lowered both MAE and RMSE. The latter approach proposed a particularly interesting "extract-and-match" paradigm, which basically totally relies on the attention mechanism of Video Transformer [111]. Additionally, ConCoNet is worth mentioning since it proposed a training strategy potentially applicable to all the other methods that brought the idea of leveraging negative exemplars to better distinguish between target and background regions. However, the model achieving state-of-the-art results is DAVE [74]. By employing a detect-and-verify strategy, DAVE combines bounding-box detection with clustering-based verification to eliminate outliers effectively. On the leaderboard, DAVE demonstrates the lowest MAE and RMSE values to date (8.66 and 32.36, respectively, on the test subset), showcasing its dominance in the reference-based paradigm. This performance leap is not solely due to architectural complexity but reflects a shift in design philosophy. Earlier regression-based methods like FamNet are vulnerable to background clutter, which can produce false density peaks and catastrophic counting errors. DAVE addresses this by decoupling the process: first generating candidate detections, then verifying them against exemplar clusters. This verification stage acts as a semantic filter, rejecting outliers and enforcing consistency with the exemplar distribution. Finally, it is worth noting that the best MAE on the validation subset is achieved by CountDiff [78]. However, this result relies on a test-time adaptation strategy, making the comparison not entirely fair since additional training occurs during inference. Without this component, performance is less competitive—MAE/RMSE of 9.15/31.97 on the validation subset and 9.97/54.89 on the test subset.
V-C Reference-less Results
In Tab. II, we report the results obtained by the surveyed reference-less approaches on FSC-147. The first reference-less approach, RepRPN-Counter [79], introduced the concept of repetitive region proposals and reached an MAE of 26.66 and an RMSE of 129.11 on the test subset. These error metrics are significantly higher than those obtained with reference-based approaches, as expected, since reference-less techniques do not utilize bounding box exemplar annotations. Subsequent methods have significantly improved performance. CounTR [81], a transformer-based reference-less framework, leverages the self-similarity of image regions and achieves competitive results, with an MAE of 14.12 and an RMSE of 108.01 on the test subset. Another notable contribution is RCC [80]. Besides achieving competitive results (an MAE of 17.12 and an RMSE of 104.53), it employs a weakly supervised learning approach that does not need point-level annotations, thereby representing a particularly interesting approach in data scarcity scenarios. In the same direction, also GCNet [82] does not use dot labels and obtained comparable results to RCC. Overall, there is no clear winner in this leaderboard. Indeed, on the one hand, the reference-less version of DAVE [74] achieves the best results in the validation split with an MAE and RMSE of 15.54 and 52.67, respectively; on the other hand, CounTR and GCNet shine in the test subset with an MAE of 14.12 and an RMSE of 102.89.
These results highlight a key trade-off in reference-less CAC: while these methods eliminate the need for exemplar annotations and offer greater flexibility, they often suffer from semantic ambiguity and background clutter. Without explicit guidance, they rely on visual repetition to infer the dominant object class, which can lead to false positives in complex scenes.
V-D Open-world Text-guided Results
In the last part of Tab. II, we show the results concerning the considered open-world text-guided approaches, which represent a recent and promising advancement in class-agnostic counting, driven by the progress of large multi-modal models. However, similar to the reference-less case, this flexibility comes at the expense of lower performance compared to reference-based techniques. The first milestone, ZSC [83], obtained an MAE and an RMSE of 22.09 and 115.17 on the test subset. Among the subsequent methods that steadily improved performance, CounTX [84] emerges as a particularly competitive approach, in some way adapting its counterpart CounTR [81] to this new setting. It achieves an impressive MAE of 15.73 and an RMSE of 106.88 on the test subset. However, the best results were contended by two open-world text-guided variants of reference-based methods—DAVE [74] and CountDiff [78]—and GroundingREC [87]. The first two excel in RMSE, indicating greater robustness to outliers, while GroundingREC achieves the best MAE (10.06 and 10.12 on the validation and test subsets, respectively).
Beyond raw performance dictated by architectural choices, the effectiveness of text-guided counting is also influenced by the richness of the textual prompts used. While most methods rely on simple class names, recent approaches have explored more structured and semantically expressive prompts. For instance, CounTX is trained on the FSC-147-D dataset, which replaces basic labels with more descriptive textual sentences. Similarly, GroundingREC is trained on the REC-8K dataset, where prompts include object attributes. Notably, both CounTX and GroundingREC rank among the top-performing models in our evaluation, suggesting that richer prompts can lead to more accurate and robust counting. These results highlight the potential of moving beyond minimal textual input toward more informative and context-aware prompting. Future research could explore automatic prompt generation, adaptive refinement mechanisms, and multimodal reasoning strategies to further improve semantic alignment and generalization in open-world scenarios, possibly also considering the contextual capabilities of large multimodal models as part of broader reasoning pipelines.
V-E Trade-offs Between Guidance and Accuracy
A central theme emerging from these results is the fundamental trade-off between performance and convenience across the three CAC paradigms. While reference-based methods consistently achieve the best results (e.g., DAVE [74] achieves an MAE of 8.66 on FSC-147 test set), they require human-annotated exemplars at inference time, limiting scalability and automation. In contrast, reference-less and open-world text-guided approaches offer greater convenience—eliminating or automating the need for exemplars—but at a measurable cost in accuracy. This “price of convenience” is starkly illustrated by the more than 70% increase in MAE observed when moving from DAVE (8.66) to its reference-less variant (15.14) or its open-world text-guided version (14.90).
However, beyond the illustrative case of DAVE, all reference-less and open-world text-guided approaches consistently show lower performance compared to reference-based techniques. Notably, the RMSE is particularly affected by the change of paradigm: the best results on the FSC-147 test set are achieved by DAVE with an RMSE of 32.36, GCNet [82] with 102.89, and the text-guided version of DAVE with 103.42, representing the top-performing models in the reference-based, reference-less, and text-guided paradigms, respectively. We argue that this large error gap in RMSE can often be attributed to the use of poor-quality or even incorrect exemplars—a challenge that becomes more pronounced in the absence of reference bounding boxes—and may result in catastrophic errors on some images. The logical conclusion is that, as of today, the core challenge of effective guidance in class-agnostic counting entails a quantifiable cost in accuracy—a trade-off that should be central to both the evaluation and future development of CAC methods.
@@ V-E Trade-offs Between Guidance and Accuracy
TABLE III: Results on the CARPK dataset [43]. We report the performance in terms of MAE and RMSE obtained with some of the methodologies discussed in this survey over the CARPK benchmark. Following previous literature, CARPK is exploited for testing cross-dataset generalization over a specific object class—vehicles. The best results are in bold.
@@
V-F Cross-dataset Generalization
Table III summarizes the results on the CARPK dataset [43] among methods that reported outcomes on this benchmark.
Fine-tuned reference-based approaches demonstrate substantial performance gains compared to non-fine-tuned ones, as expected, since the latter do not see the vehicle category at all. In contrast, experiments on CARPK with fine-tuned models diverge slightly from the CAC paradigm, as the model is tested on the same class seen during fine-tuning. Rather than evaluating generalization to unseen categories, as in the standard CAC setting, this scenario assesses the ability to learn from a small number of exemplars and generalize across a dataset containing objects of the same category. Specifically, CACViT [73] leads the pack of fine-tuned techniques (MAE: 4.91, RMSE: 6.49). Its performance likely stems from the use of a pre-trained Vision Transformer that performs unified feature extraction and matching via attention mechanisms. The model processes the query image and exemplar patches as token sequences—referred to as query and exemplar tokens—which are concatenated and jointly fed into the transformer encoder. Within this unified token space, self-attention extracts contextual features from both inputs and simultaneously performs similarity matching between them. This design appears particularly effective in learning from very few exemplars (12 in this case), and demonstrates robustness in generalizing across the dataset. BMNet+ [63] and SAFECount [66] also deliver strong results (MAEs of 5.76 and 5.33, respectively), leveraging advanced similarity modules and data-driven refinement strategies that enhance adaptability.
Non-fine-tuned methods generally show higher error rates, yet recent models like LOCA [67], SATCount [70], CACViT [73], and SSD [69] narrow the gap with fine-tuned approaches. This set of experiments fully reflects the CAC setting, since models have not seen the vehicle category during training. The goal here is to evaluate their ability to count a specific object category in a new dataset under domain shift. Notably, UPC [76] achieves the best performance among non-fine-tuned methods (MAE: 7.84, RMSE: 9.74), and remains competitive even in open-world text-guided settings (MAE: 7.62, RMSE: 9.71). UPC’s strong generalization may be attributed to its iterative refinement mechanism, which progressively updates the predicted density map by reusing it as a new prompt. This feedback loop allows the model to converge toward consistent outputs and may help suppress spurious activations in unfamiliar domains. Additionally, the use of a contrastive training strategy—where the model learns to distinguish between prompts that match the target object and those that do not—encourages the network to focus on semantically meaningful regions. These mechanisms likely improve the model’s robustness to domain shift by reinforcing object-specific consistency and reducing reliance on dataset-specific context. Notably, VLCounter [86] stands out among open-world text-guided methods (MAE: 6.46, RMSE: 8.68), but UPC’s performance remains notable given its unified prompt representation and refinement strategy.
V-G Textual Semantic Understanding
@@
Recently, [126] highlighted notable weaknesses in state-of-the-art open-world text-guided CAC methods. Specifically, the authors argued that these models often fail to accurately identify which object class needs to be counted based on the textual description. Instead, they tend to count instances of the predominant class, disregarding the true intent of the prompt. To address these shortcomings, the authors introduced a novel benchmark, including a test suite specifically designed to evaluate these limitations in current evaluation systems.
@@
Table IV presents results on the FSC-147 dataset, focusing on their negative-label test. This test evaluates the performance of the model on single-class images by prompting it with textual descriptions referencing absent object classes. [126] introduced also two metrics for the assessment of this test: (i) the normalized mean of negative predictions (NMN), which is the absolute counting error computed by prompting the model with the negative classes, normalized by the ground truth of the positive class, and (ii) the positive class count nearness (PCCN) that mixes positive and negative class predictions, providing an overall quantitative assessment of strong failures in the model. More details can be found in [126]. DAVE [74] performs exceptionally well across nearly all metrics proposed in [126], standing out as the best-performing model in this challenging scenario. Although the authors noted that DAVE occasionally produces catastrophic errors due to sporadic outliers, it remains the top performer overall in this evaluation.
V-H Computational Complexity
Table V reports the computational complexity—measured in GFLOPs and number of parameters—of the most representative CAC methods, i.e., those highlighted in Fig. 3. Specifically, we include values reported in the original papers when available, and compute them for methods that provide public code but omit complexity metrics. Models lacking both reported values and code are excluded from the analysis.
TABLE V: Computational complexity analysis. We report the computational complexity of the most representative CAC methods in terms of GFLOPs and number of parameters. “-” indicates that the values were not computed because the authors did not provide a pretrained model. The best results are highlighted in bold.
Method GFLOPs $\downarrow$ #Parameters $\downarrow$
Reference-based
FamNet [32] 55 26M
RCAC [62] - 10M
BMNet+ [63] 27 13M
SAFECount [66] 366 32M
LOCA [67] 80 37M
SSD [69] - 36M
CACViT [73] 89 99M
DAVE [74] 1,515 47M
TFPOC [75] 102 94M
CounTR\$ [81] 84 99M
Open-world Text-guided
ZSC [83] - 12M
CounTX [84] 50 161M
CLIP-Count [85] 27 166M
VLCounter [86] 68 1.4M
GroundingREC [87] 69 173M
VA-Count [88] - 99M
$\ddagger$ computed in this work; \$ reference-based modification of a reference-less method.
The results reveal substantial variability in computational requirements across paradigms and architectures. Among reference-based methods, models such as FamNet [32] and BMNet+ [63] are relatively lightweight, requiring 55 and 27 GFLOPs, respectively, with moderate parameter counts (26M and 13M). In contrast, more recent architectures like SAFECount [66] and LOCA [67] show a marked increase in both GFLOPs (366 and 80) and parameters (32M and 37M), reflecting the adoption of deeper backbones and more sophisticated matching modules. DAVE [74] represents a particularly interesting case since it stands out as the state-of-the-art in terms of accuracy, achieving the lowest MAE and RMSE on both FSC-147 [32] and CARPK [43]. However, this performance comes at a very high computational cost: 1515 GFLOPs and 47M parameters—an order of magnitude higher than most other methods. This makes DAVE ideal for high-resource environments, but potentially unsuitable for edge deployment.
In the open-world text-guided paradigm, computational demands are generally higher due to the integration of large vision-language models. For instance, CounTX [84] and GroundingREC [87] reach parameter counts of 161M and 173M, respectively. Notably, VLCounter [86] achieves a remarkably low parameter count (1.4M) while maintaining competitive GFLOPs (68), thanks to efficient prompt-tuning and segment-aware skip connections.
Overall, the analysis reveals that higher computational complexity does not always correlate with better performance. While DAVE justifies its cost with top-tier accuracy, other models like CACViT, SSD, and BMNet+ offer a more favorable trade-off between efficiency and accuracy. These findings suggest that architectural innovations—rather than brute-force scaling—are key to achieving practical and scalable CAC solutions.
Figure 6: Qualitative results of reference-based CAC methods. We report some qualitative results of the most representative reference-based CAC techniques. On the left, we show three easy samples, while on the right, three challenging samples that result in large counting errors.
Figure 7: Qualitative results of open-world text-guided CAC methods. We report some qualitative results of the most representative open-world text-guided CAC techniques. On the left are three easy samples, while on the right are three challenging samples that result in large counting errors.
V-I Qualitative Comparison
Fig. 6 presents qualitative results for reference-based CAC techniques. We compare the outputs of the most representative CAC models (see Fig. 3) for which pretrained weights are publicly available. Specifically, we show both easy and hard samples, with the latter often corresponding to failure cases. Note that, for TFPOC [75], density maps were generated from the predicted masks by placing a small circle at the centroid of each mask.
Overall, DAVE [74] confirms its state-of-the-art performance. However, it is worth noting that even DAVE, like all other models, struggles with certain challenging samples. One recurring challenge appears to be related to object density—objects tend to overlap in crowded scenes. This issue becomes more pronounced when additional elements such as background clutter and distractors are present, as in the second hard sample. In contrast, it is less evident in scenarios where only the target objects are present and are relatively uniformly distributed, as in the third hard sample. A particularly interesting failure case is illustrated by the first hard sample. Despite the absence of occlusion and the seemingly simple scene, all models consistently fail by counting the lenses of the glasses rather than the glasses as whole objects. This suggests that the models are unable to build effective representations of the exemplars and instead rely on detecting repetitive patterns in the image.
Fig. 7 presents instead qualitative results for the most representative open-world text-guided CAC methods. Note that, for GroundingREC [87], density maps were generated by placing a small circle at the centroids of the generated bounding boxes. The most pronounced and interesting failure case is observed in the second hard sample. As previously noted by [126] and discussed in Sec. V-G, this error stems from the limited ability of the models to understand the semantic content of textual descriptions. In particular, they tend to count repeated objects regardless of the actual meaning conveyed by the text. This issue is clearly illustrated in the image, which contains a large number of objects from two distinct categories. Only DAVE [74], and to a lesser extent GroundingREC [87], can successfully address this challenge.
@@ VI Conclusion and Open Challenges
Class-agnostic counting has become a pivotal field in object counting, addressing the limitations of traditional class-specific methods. Unlike conventional approaches that require extensive labeled data for each object category, CAC enables models to count objects across arbitrary categories, including those unseen during training. For the first time, this survey reviewed the evolution of CAC methodologies, categorizing them into three paradigms: reference-based, reference-less, and open-world text-guided approaches. Each paradigm brings distinct advantages and trade-offs, reflecting the diverse demands of real-world applications. Reference-based methods, which rely on annotated exemplars, set the benchmark for accuracy and robustness. However, their dependence on labeled exemplars limits scalability and automation. Reference-less approaches address this by automatically identifying dominant object classes, offering greater flexibility but still lagging behind in performance. Open-world text-guided methods further increase adaptability by allowing object specification via textual prompts, yet—like reference-less methods—this flexibility comes at the cost of reduced accuracy compared to exemplar-based approaches.
Through an in-depth analysis of state-of-the-art methods on benchmarks such as FSC-147 [32] and CARPK [43], this survey highlights both notable advancements and persistent challenges. While DAVE [74] leads the reference-based paradigm on FSC-147 (MAE: 8.66, RMSE: 32.36), the reference-less and text-guided leaderboards remain less conclusive. CounTR [81] and GCNet [82] stand out among reference-less methods, the latter notable for its weak supervision. In the text-guided setting, GroundingREC [87] achieves the lowest MAE, while DAVE remains robust to outliers. Cross-dataset evaluations on CARPK reveal that CACViT [73], UPC [76], and VLCounter [86] generalize best across paradigms. Qualitative results further confirm that even top-performing models can fail in cluttered or ambiguous scenes, often due to poor or missing guidance. Moreover, recent benchmarks [126] show that open-world models frequently misinterpret prompts, highlighting limitations in semantic understanding.
@@
Open challenges span all paradigms. A major limitation is the lack of diverse and publicly available datasets. The gold standard, FSC-147, has several drawbacks. First, it contains a limited number of images and object classes. Furthermore, as pointed out in [126], almost all images contain only a single class of objects. Multi-class images would be a desirable feature for new datasets, as objects belonging to classes different from the one being counted could act as distractors for the counting approach. The lack of diverse and large-scale datasets also leads to challenges in designing learning approaches suited for data-scarce scenarios. Amnog the analyzed methods, only RCC [80] and GCNet [82] operate with weak supervision, avoiding the use of dot annotations. This setting is particularly promising, as it could help mitigate the challenge of creating large-scale labeled datasets for the counting task. Recently, an emerging direction is to incorporate completely unsupervised components into CAC pipelines [127, 128], although current performance remains far from state-of-the-art.
Other challenges are more specific to open-world text-guided methodologies, which appear to represent the future. These models have been shown to occasionally misinterpret the meaning of prompts. Although some architectures demonstrate relative robustness, further improvements are needed. New architectural designs and learning strategies—such as contrastive learning over classes not present in the images—should be explored. Another particularly interesting direction concerns the use of richer textual inputs to enable context-aware prompting. Future research could investigate automatic prompt generation techniques to produce more informative and targeted textual queries. While current large multimodal models are not yet effective for counting tasks, partly due to limitations in visual token capacity, their contextual priors and exposure to diverse visual-textual data suggest potential for future exploration, especially as architectures evolve to better handle fine-grained visual details. Beyond these technical challenges, our analysis underscores a broader issue: the convenience offered by reference-less and text-guided paradigms comes at a measurable cost in accuracy. Bridging this gap—by developing methods that match the guidance quality of human-provided exemplars without manual intervention—remains a key open challenge. Finally, our computational complexity analysis reveals that top-performing models such as DAVE achieve state-of-the-art accuracy at the cost of significantly higher GFLOPs and parameter counts. This highlights a critical trade-off between performance and efficiency, which must be carefully considered when designing CAC systems for real-world deployment, particularly in resource-constrained environments.